Shirley bougth 4 jackets for $ 90.30 if each jacket cost $21 before tax was added what tax rate did she pay
Answer by DrBeeee(684)

The amount paid for tax, T, is
(1) T = 93.3 - 4*21 or
(2) T = 9.3
The tax rate is applied the the original cost of $84, giving us the rate, r, as
(3) r = 9.3/84 or
(4) r = 11.07%
Is (84 + (11.07/100)*64 = 93.3)?
Is (84 + 9.3 = 93.3)?
Is (93.3 = 93.3)? Yes
Answer: Shirley paid a 11.07% tax rate.